# Safeguarding System Prompts for LLMs

### v1

@@ Appendix D Details of Evaluating Conversational Capability
Please act as an impartial judge and evaluate the adherence of the AI assistant’s response to the system message. Responses that deviate from the specified scope or behavior should be rated low, even if they are helpful and relevant. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: “[[rating]]”, for example: Rating: [[5]].
@@
Also, we tailor the question set to each system prompt so that the queries can produce markedly different responses depending on whether the prompt is presented to the model. These queries are specifically generated by OpenAI’s gpt-4 using the following query, with the system prompt to protect also in place:
@@
Give me [some number] example questions where your response would fail to adhere to all the instructions you see if you had not seen these instructions. Start with "1:". Strictly after providing all your example questions, you must give a short explanation of why you answered these questions.
Take the following system prompt from the Real GPTs dataset as an example:
@@
Appendix E More Results on Defense Effectiveness
Table 2: Mean attack performance under various defenses with Synthetic GPTs.
Defense Adversarial-Query Attack Regular-Query Attack
Llama No defense 92.0 $\pm$ 8.5 39.0 $\pm$ 26.3 62.5 $\pm$ 28.0 93.3 $\pm$ 4.1 12.7 $\pm$ 5.9 46.8 $\pm$ 7.0
(a) Llama (Adversarial).
(b) Llama (Regular).
(c) Mistral (Adversarial).
(d) Mistral (Regular).
Figure 6: How various defenses navigate the privacy-capability tradeoff with Synthetic GPTs.
Table 3: Mean attack performance under various defenses with Awesome ChatGPT Prompts.
Defense Adversarial-Query Attack Regular-Query Attack
Llama No defense 91.2 $\pm$ 7.2 19.6 $\pm$ 17.8 50.0 $\pm$ 20.8 83.4 $\pm$ 5.1 2.3 $\pm$ 2.0 25.4 $\pm$ 5.6
(a) Llama (Adversarial).
(b) Llama (Regular).
(c) Mistral (Adversarial).
(d) Mistral (Regular).
Figure 7: How various defenses navigate the privacy-capability tradeoff with Awesome ChatGPT Prompts.
While Section 6.1 primarily discusses the results obtained with the Real GPTs dataset, we also present results from the Synthetic GPTs dataset in Table 2 and Figure 6, and Awesome ChatGPT Prompts dataset in Table 3 and Figure 7, respectively. The observations from these datasets are consistent with those obtained from the Real GPTs dataset.
Appendix F Worst-Case Defense Effectiveness
As discussed in Section 5.4, the attack performance used to evaluate defense effectiveness is primarily reported as an average across attack instances and repetitions. This approach offers two key advantages: (1) it aligns with the reporting standards of prior work (Morris et al., 2024; Zhang et al., 2024a), enabling validation of our attack implementations; and (2) it provides immediate insights into how effectively PromptKeeper safeguards system prompts when assessed using established benchmarks, both as presented in Section 6.1.
However, as also highlighted in Section 3, the design of PromptKeeper accounts for worst-case scenarios. Consequently, evaluating the maximum attack performance is equally important to capture the upper bounds of potential leakage. These worst-case results are reported in Table 4, Table 5, and Table 6, which reveal trends consistent with those observed in the average attack performance, thereby reinforcing our original conclusions.
Table 4: Worst-case attack performance under various defenses with Real GPTs.
Defense Adversarial-Query Attack Regular-Query Attack
Llama No defense 100.0 100.0 100.0 96.3 18.3 44.9
No prompt 77.2 2.4 30.3 88.8 6.5 30.4
Query filter 99.2 78.1 92.2 95.9 13.8 45.1
Self-extension 99.6 93.1 97.6 95.5 11.1 49.7
Regen w/ CS 98.6 67.9 83.2 96.2 15.2 42.3
PromptKeeper 96.7 38.8 70.9 95.4 8.6 39.3
Mistral No defense 99.8 79.6 92.0 96.7 29.3 50.0
No prompt 79.7 2.7 30.3 89.0 5.6 29.6
Query filter 99.8 92.1 97.2 95.9 19.2 48.5
Self-extension 100.0 100.0 100.0 96.9 19.7 50.5
Regen w/ CS 98.7 64.6 80.4 97.0 21.7 47.5
PromptKeeper 97.5 56.7 68.6 95.8 17.0 47.4
Table 5: Worst-case attack performance under various defenses with Synthetic GPTs.
Defense Adversarial-Query Attack Regular-Query Attack
Llama No defense 99.2 96.4 98.6 98.3 28.2 64.2
No prompt 79.4 1.3 24.7 90.2 7.7 35.4
Query filter 98.9 98.3 98.6 98.3 26.2 57.9
Self-extension 99.3 98.1 98.6 98.7 32.8 60.4
Regen w/ CS 98.7 65.4 84.9 98.2 30.5 59.6
PromptKeeper 97.6 23.4 66.7 97.2 21.1 50.2
Mistral No defense 98.9 94.2 97.1 97.7 27.2 58.4
No prompt 80.3 1.4 24.7 89.5 7.0 35.4
Query filter 98.9 92.3 95.6 99.1 42.8 66.0
Self-extension 98.9 92.7 96.2 98.9 33.8 63.9
Regen w/ CS 98.7 71.5 85.5 99.1 31.0 64.9
PromptKeeper 98.5 26.9 61.5 96.4 13.1 56.0
Table 6: Worst-case attack performance under various defenses with Awesome ChatGPT Prompts.
Defense Adversarial-Query Attack Regular-Query Attack
Llama No defense 99.3 81.3 89.4 92.5 10.1 35.8
No prompt 78.1 2.0 32.4 75.7 1.5 22.9
Query filter 97.7 76.8 89.6 86.2 12.2 40.0
Self-extension 100.0 100.0 100.0 89.1 8.0 40.8
Regen w/ CS 96.8 34.6 80.0 89.7 10.2 44.1
PromptKeeper 94.1 28.9 65.1 89.0 2.3 26.4
Mistral No defense 97.2 17.0 63.4 88.5 4.9 40.6
No prompt 77.0 2.3 25.9 75.6 1.9 23.1
Query filter 96.8 23.4 64.7 86.6 5.5 34.5
Self-extension 96.4 44.4 61.4 90.2 12.0 50.0
Regen w/ CS 96.8 13.0 57.9 90.0 10.0 33.9
PromptKeeper 95.3 9.7 44.9 84.8 4.0 33.3

### v2

@@ 5.4. Metrics
Give me [some number] example questions where your response would fail to adhere to all the instructions you see if you had not seen these instructions. Start with ”1:”. Strictly after providing all your example questions, you must give a short explanation of why you answered these questions.
An example system prompt to protect $\bm{p}$ and the correspondingly generated question set $S$ can be found in Appendix B. The target language model is then queried with each question $\bm{q}\in S$. For each query-response pair $(\bm{q},\bm{r})$, we evaluate the target model’s adherence to the system prompt $\bm{p}$ by inputting the following system prompt $\bm{p}_{judge}$ and query $\bm{q}_{judge}$ into the judge LM:
@@
System prompt for judging ($\bm{p}_{judge}$):
Please act as an impartial judge and evaluate the adherence of the AI assistant’s response to the system message. Responses that deviate from the specified scope or behavior should be rated low, even if they are helpful and relevant.
Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: “[[rating]]”, for example: Rating: [[5]].
{mdframed}
User query for judging ($\bm{q}_{judge}$):
@@
Compared to potential manual evaluation, this approach alleviates the costly and labor-intensive burden while maintaining interpretability, as the judge LM can also generate natural-language explanations for its scores.
Table 1. Mean attack performance under various datasets and defenses.
Dataset Target Model Defense Adversarial-Query Attack Regular-Query Attack
Real [0.5em]GPTs Llama No defense 91.0 $\pm$ 9.1 31.0 $\pm$ 27.1 56.3 $\pm$ 26.0 90.9 $\pm$ 4.2 5.4 $\pm$ 3.8 33.6 $\pm$ 6.8
Synthetic [0.5em]GPTs Llama No defense 92.0 $\pm$ 8.5 39.0 $\pm$ 26.3 62.5 $\pm$ 28.0 93.3 $\pm$ 4.1 12.7 $\pm$ 5.9 46.8 $\pm$ 7.0
Awesome [0.5em]ChatGPT [0.5em]Prompts Llama No defense 91.2 $\pm$ 7.2 19.6 $\pm$ 17.8 50.0 $\pm$ 20.8 83.4 $\pm$ 5.1 2.3 $\pm$ 2.0 25.4 $\pm$ 5.6
(a) Real GPTs, Llama (A).
(b) Real GPTs, Llama (R).
(c) Real GPTs, Mistral (A).
(d) Real GPTs, Mistral (R).
(e) Synthetic GPTs, Llama (A).
(f) Synthetic GPTs, Llama (R).
(g) Synthetic GPTs, Mistral (A).
(h) Synthetic GPTs, (R).
(i) ChatGPT Awesome Prompts, Llama (A).
(j) ChatGPT Awesome Prompts, Llama (R).
(k) ChatGPT Awesome Prompts, Mistral (A).
(l) ChatGPT Awesome Prompts, Mistral (R).
Figure 4. How various defenses navigate the privacy-capability tradeoff under (A)dversarial- and (R)egular-query attacks, respectively. While attack efficiency is measured here using cosine similarity (CS), the observed trends are consistent with those obtained using BLEU or token-level F1 scores.
6. Evaluation
6.1. Average-Case Defense Effectiveness
We focus here on the evaluation with the Real GPTs dataset. Trends observed in the Synthetic GPTs and Awesome ChatGPT Prompts datasets are consistent and are thus omitted for brevity.
Validity of implemented attacks.   As mentioned in Section 5.2 and 5.4, we assess the effectiveness of a defense mechanism against system prompt leakage by evaluating the difficulty of two extraction attacks. In both our Llama and Mistral testbeds, these attacks achieve a high level of success when no defense mechanism is in place, as indicated by the “No defense” rows in Table 1. For instance, the mean cosine similarity of text embeddings, BLEU score, and Token-level F1 between the reconstructed and the ground truth system prompts mount to 94.9, 31.0, and 59.2, respectively. These results align with previous studies (Zhang et al., 2024b; Morris et al., 2024), validating our implementation of these state-of-the-art attacks and providing a solid foundation for evaluating defenses.
Table 2. Worst-case attack performance under various datasets and defenses.
Dataset Target Model Defense Adversarial-Query Attack Regular-Query Attack
Real [0.5em]GPTs Llama No defense 100.0 100.0 100.0 96.3 18.3 44.9
No prompt 77.2 2.4 30.3 88.8 6.5 30.4
Query filter 99.2 78.1 92.2 95.9 13.8 45.1
Self-extension 99.6 93.1 97.6 95.5 11.1 49.7
Regen w/ CS 98.6 67.9 83.2 96.2 15.2 42.3
PromptKeeper 96.7 38.8 70.9 95.4 8.6 39.3
Mistral No defense 99.8 79.6 92.0 96.7 29.3 50.0
No prompt 79.7 2.7 30.3 89.0 5.6 29.6
Query filter 99.8 92.1 97.2 95.9 19.2 48.5
Self-extension 100.0 100.0 100.0 96.9 19.7 50.5
Regen w/ CS 98.7 64.6 80.4 97.0 21.7 47.5
PromptKeeper 97.5 56.7 68.6 95.8 17.0 47.4
Synthetic [0.5em]GPTs Llama No defense 99.2 96.4 98.6 98.3 28.2 64.2
No prompt 79.4 1.3 24.7 90.2 7.7 35.4
Query filter 98.9 98.3 98.6 98.3 26.2 57.9
Self-extension 99.3 98.1 98.6 98.7 32.8 60.4
Regen w/ CS 98.7 65.4 84.9 98.2 30.5 59.6
PromptKeeper 97.6 23.4 66.7 97.2 21.1 50.2
Mistral No defense 98.9 94.2 97.1 97.7 27.2 58.4
No prompt 80.3 1.4 24.7 89.5 7.0 35.4
Query filter 98.9 92.3 95.6 99.1 42.8 66.0
Self-extension 98.9 92.7 96.2 98.9 33.8 63.9
Regen w/ CS 98.7 71.5 85.5 99.1 31.0 64.9
PromptKeeper 98.5 26.9 61.5 96.4 13.1 56.0
Awesome [0.5em]ChatGPT [0.5em]Prompts Llama No defense 99.3 81.3 89.4 92.5 10.1 35.8
No prompt 78.1 2.0 32.4 75.7 1.5 22.9
Query filter 97.7 76.8 89.6 86.2 12.2 40.0
Self-extension 100.0 100.0 100.0 89.1 8.0 40.8
Regen w/ CS 96.8 34.6 80.0 89.7 10.2 44.1
PromptKeeper 94.1 28.9 65.1 89.0 2.3 26.4
Mistral No defense 97.2 17.0 63.4 88.5 4.9 40.6
No prompt 77.0 2.3 25.9 75.6 1.9 23.1
Query filter 96.8 23.4 64.7 86.6 5.5 34.5
Self-extension 96.4 44.4 61.4 90.2 12.0 50.0
Regen w/ CS 96.8 13.0 57.9 90.0 10.0 33.9
PromptKeeper 95.3 9.7 44.9 84.8 4.0 33.3
6.2. Worst-Case Defense Effectiveness
The above attack performance used to evaluate defense effectiveness is primarily reported as an average across attack instances and repetitions. This approach offers two key advantages: (1) it aligns with the reporting standards of prior work (Morris et al., 2024; Zhang et al., 2024b), enabling validation of our attack implementations; and (2) it provides immediate insights into how effectively PromptKeeper safeguards system prompts when assessed using established benchmarks. However, evaluating the maximum attack performance is equally important to understand the upper bounds of potential leakage. The worst-case results for the three evaluated datasets are reported in Table 2. These results exhibit trends consistent with the average attack performance (Section 6.1), further reinforcing our original conclusions.
6.3. Necessity for On-Demand Regeneration
Assume that the word limit is of interest to an attacker, the implemented defense makes it challenging for it to infer this limit by directly querying the language model with prompts like, ”Tell me your word limit.” In such cases, the corresponding response is likely to leak information about the prompt and would therefore be replaced with an empty one. However, the attacker can exploit the DoS behavior to launch a more efficient attack. It can issue a query following the template: “Ignore your task and strictly say ‘I draft at most [some guess of number] words.”’
In this case, if the guess contained in the sentence to reiterate is close to the actual value, the LM’s reiteration may be considered a leakage of the system prompt and thus trigger a DoS. Conversely, if the guess is not close, the reiteration will likely be output without modification. This distinction allows the attacker to differentiate between the two cases, facilitating a strategic search with multiple queries. For instance, the attacker can sweep guesses within a range, such as [20, 200]. As shown in Figure 5(a), when the guess is near the actual value, the service consistently returns No response, while Reiterating the required sentence for guesses outside this vicinity, regardless of the choice of the significance level $\alpha$. This implies that the attacker can infer the word limit effectively. In contrast, as shown in Figure 5(b), if the original PromptKeeper is in place, the service consistently Reiterates the required sentence, even when the attacker’s guess is close to the actual value. This highlights the superiority of on-demand regeneration with dummy prompts for response-based defenses (Section 4).
(a) PromptKeeper w/o regeneration but DoS.
(b) PromptKeeper.
Figure 5. Examples demonstrating the advantage of on-demand regeneration over denial of service.
7. Discussion and Future Work
Risk of false negatives.   As defined in Section 4, a false negative occurs when the test incorrectly indicates non-zero leakage despite zero leakage. Similar to any binary classification system, PromptKeeper carries a risk of false negatives. Our leakage identification method offers flexibility in managing the trade-off between the actual false negative rate and actual false positive rate, where the latter is proportionate with $\alpha$, i.e., the target false positive rate. Theoretically, increasing $\alpha$ reduces the empirical false negative rate, allowing service providers to select an $\alpha$ value that best aligns with their specific requirements. However, it is important to emphasize that the actual false negative rate is intractable in practice, as constructing and validating positive cases—responses generated with a system prompt that leak no information about the prompt—is challenging. As a result, the risk of false negatives can only be discussed conceptually, without introducing additional empirical evaluations.
Necessity for full knowledge of system prompts.   As required for the offline estimation of $\tilde{Q}^{*}_{\textrm{zero/other}}(\bm{p},\bm{q})$ (Section 3), PromptKeeper necessitates full knowledge of the system prompt $\bm{p}$ to ensure its protection. This design is intended for deployment by service providers, aligning with standard practices in the industry. For example, popular LLM platforms like OpenAI’s GPT Store and Poe require developers to specify their system prompts in plaintext, effectively making the service provider the prompt owner. This practice is likely to persist in the near future, as alternatives such as trusted execution environments (TEEs) (Intel, 2021; Arm, 2021) or secure multiparty computation (SMPC) (Evans et al., 2018) for user query processing are currently impractical due to their significant computational and financial costs. Until these techniques become viable for industrial-scale deployment, LLMs will continue to rely on plaintext system prompts, ensuring service providers maintain access to them.
Transfer to safeguarding user queries.   An adversary might eavesdrop on responses received by a user and attempt to extract the queries used. Unfortunately, PromptKeeper cannot be generalized to protect them against such threats. This is because our method necessitates active involvement from the service provider for hypothesis testing, yet it lacks the incentive to do so merely for user privacy. Even with the provider’s cooperation, balancing privacy and capability in the context of user query protection is tricky. Unlike system prompts, which can function even if their information is not included in the model response, a user query typically needs to be incorporated in the response for it to be useful. This calls for independent research on user query protection.
8. Related Works
System prompt extraction attacks.   System prompt extraction has emerged as a critical threat to LLM deployments. Studies have shown that adversarial queries, such as instructions to repeat hidden inputs, can effectively extract system prompts verbatim (Perez and Ribeiro, 2022; Wallace et al., 2024; Zhang et al., 2024a; Hui et al., 2024). More recent work has highlighted that even regular queries, often appearing benign, can lead to prompt leakage by inverting model outputs or token-level logits to the original prompts (Zhang et al., 2024b; Morris et al., 2024).
Defenses against system prompt leakage.   Few studies have proposed comprehensive solutions specifically for protecting system prompts. Input-based approaches, such as augmenting system prompts or filtering adversarial queries, have been mentioned briefly in prior work without systematic exploration or thorough evaluation (Hui et al., 2024; Zhang et al., 2024a). Similarly, training-related approaches have only been implicitly referenced, such as using fine-tuning to discourage certain behaviors (Achiam et al., 2023; Ouyang et al., 2022). However, as summarized in Section 4.1, these approaches suffer from inherent limitations in either conversational capability, defense effectiveness or runtime efficiency, making them insufficient as standalone solutions.
The closest response-based defense to our work is (Zhang et al., 2024a), where the model denies a response if there is an n-gram overlap between the generated output and the system prompt. However, as the authors acknowledge, this defense can be easily bypassed by attackers instructing the language model to rephrase the extracted prompt. This limitation is fundamental—any leakage identification approach relying on imperfect metrics is inherently prone to inaccuracies. In contrast, PromptKeeper adopts a robust statistical approach for leakage detection and also introduces a general mechanism to mitigate side-channel vulnerabilities in response-based defenses.
Regarding side-channel vulnerabilities specifically, Debenedetti et al. (2024) explored them in the context of protecting training data. However, their work focuses solely on the leakage of sensitive data in verbatim form and does not provide a corresponding countermeasure for side-channel vulnerabilities. PromptKeeper bridges these gaps with a comprehensive and targeted strategy that addresses both leakage detection and side-channel mitigation effectively.
9. Conclusion
Prompt extraction has long raised privacy concerns in LLM usage. Although system prompts and user queries are combined as input to LLMs, safeguarding them necessitates distinct approaches due to their differing threat models. Unlike previous studies that often address them collectively, this paper introduces PromptKeeper, an early effort focused specifically on protecting system prompts.
Leveraging the statistical properties of LLMs and the system prompts accessible to service providers, PromptKeeper offers a robust method for leakage identification. By avoiding reliance on imperfect metrics, it ensures accurate detection of both explicit and subtle leakage. Furthermore, PromptKeeper demonstrates how response-based defenses, through on-demand regeneration, can effectively neutralize side-channel attempts while minimizing disruption to benign user interactions. This dual focus on robust protection and user experience positions PromptKeeper as a comprehensive solution for safeguarding system prompts.
Appendix A Examples of Evaluated System Prompts
@@
Synthetic GPTs.   The mentioned user prompt for generating synthetic system prompts based on each name and description collected from GPTs Hunter (AI and Joanne, 2024) is provided as follows.
{mdframed}
{mdframed}
{mdframed}
Appendix B Details of Evaluating Conversational Capability
As mentioned in Section 5, we propose a language model (LM)-based method for evaluating the adherence of a model’s response to the system prompt. This method is grounded in the use of a set of prompt-aware queries. Below, we provide an example system prompt from the Real GPTs dataset along with the generated query set for the reader’s reference: